### v2

# Zenith: Scaling up Ranking Models for Billion-scale Livestreaming Recommendation

Abstract.
Accurately capturing feature interactions is essential in recommender systems, and recent trends show that scaling up model capacity could be a key driver for next-level predictive performance. While prior work has explored various model architectures to capture multi-granularity feature interactions, relatively little attention has been paid to efficient feature handling and scaling model capacity without incurring excessive inference latency. In this paper, we address this by presenting Zenith, a scalable and efficient ranking architecture that learns complex feature interactions with minimal runtime overhead. Zenith is designed to handle a few high-dimensional Prime Tokens with Token Fusion and Token Boost modules, which exhibits superior scaling laws compared to other state-of-the-art ranking methods, thanks to its improved token heterogeneity. Its real-world effectiveness is demonstrated by deploying the architecture to TikTok Live, a leading online livestreaming platform that attracts billions of users globally. Our A/B test shows that Zenith achieves +1.05%/-1.10% in online CTR AUC and Logloss, and realizes +9.93% gains in Quality Watch Session / User and +8.11% in Quality Watch Duration / User.
Ranking, Scaling Laws, Feature Interaction, Recommender Systems
* Equal Contributions.
$\dagger$ Work done during internship at TikTok.
$\ddagger$ Corresponding author.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systemsccs: Information systems Learning to rankccs: Computing methodologies Neural networksccs: Information systems Multimedia streaming
1. Introduction
Recommender systems play a vital role in modern online platforms, helping users discover relevant content, products, or services. Feature interaction has been recognized as a key mechanism in recommendation models. In recent years, deep learning has emerged as a powerful paradigm for modeling such interactions (Naumov et al., 2019; Wang et al., 2021, 2017; Zhang et al., 2023; Xu et al., 2023). By learning hierarchical, non-linear representations of users, items, and other context information, deep neural networks can capture complex feature interactions and high-order correlations.
With the availability of increasingly large-scale user behavior datasets and the demonstrated effectiveness of scaling laws in large language models (Kaplan et al., 2020), recent research has also focused on scaling recommendation models. Some efforts target the expansion of sparse feature embeddings (Mudigere et al., 2021; Lian et al., 2022), while others concentrate on enhancing dense model components or sequence modeling layers above the embedding (Ardalani et al., 2022; Zhang et al., 2023).
Despite these advancements, deep learning-based recommendation systems still face critical challenges as model size increases. Firstly, many neural architectures, such as transformers, are originally designed for tasks like natural language processing. While efficient in their native domains, they are not directly optimized for the unique characteristics of the recommendation setting—particularly the increasingly large and diverse sparse feature embeddings that require careful token-level handling. Hiformer (Gui et al., 2023) has shown that the vanilla attention mechanism is not expressive enough for ranking models, but more work is needed to gain further insight. Secondly, most existing work does not provide flexible designs for scaling up. Instead, they can only scale up the number of layers or the hidden units inside (Wang et al., 2021; Zhang et al., 2023; Xu et al., 2023). This brute-force scaling method usually leads to diminishing returns and complicates deployment across different feature sets and frameworks. Thirdly, real-world large-scale online recommender systems need to handle millions of requests simultaneously while maintaining latency at millisecond levels, making it difficult to retain the benefits of complex architectures without incurring prohibitive computational costs for traditional models.
To address these challenges, we propose Zenith (Optimized Efficient Neural Interaction with Tokenwise Handling), a novel ranking model architecture that enriches model expressiveness without compromising online performance. The key rationale behind Zenith is to treat a few, high-dimensional feature tokens as fundamental processing units. Specifically, our starting embedding layer projects the input sparse features into separate embeddings and combines feature embeddings within the same category (e.g., user demographic features, user past interaction sequences) into higher-dimensional Prime Tokens with a significantly reduced token count, compared to the number of raw features. To handle these Prime Tokens, we design two dedicated components: Token Fusion and Token Boost modules. The Token Fusion module mirrors a cross network by modeling pairwise interactions across the input tokens, while the Token Boost module processes each token independently, preserving and refining its individual semantics. As we shall demonstrate in this paper, the independent handling of each Prime Token is crucial in maintaining token heterogeneity across interaction layers for better scaling up potential, while significantly reducing the computational overhead. The superiority of Zenith to other state-of-the-art ranking architectures at various model scales is demonstrated via offline and online evaluations through our TikTok Live recommendation platform, which serves billions of users globally.
To summarize, the main contributions of this paper are:
  • We propose Zenith, a novel ranking model architecture that features Prime Token design and tokenwise handling. We provide insights on the importance of token heterogeneity in scaling up ranking models and showcase how our design meticulously strengthens this important property.
  • We introduce two effective designs for Token Fusion and Token Boost modules in Zenith respectively. Specifically, we propose Retokenized Self-Attention (RSA) and Tokenwise Multi-Head Self-Attention (TMHSA) for Token Fusion and Tokenwise SwiGLU (TSwiGLU) and Tokenwise Sparse Mixture of Experts (TSMoE) for Token Boost. These designs exhibits superior scaling laws with respect to both model parameters and compute compared to state-of-the-art ranking architectures.
  • We present offline evaluations and online A/B test results based on billions of real-world user interaction data from the TikTok Live recommendation platform. Compared to the baseline, the proposed new architecture achieves +1.05%/-1.10% in offline AUC/logloss and +9.93% in Quality Watch Session / User and +8.11% in Quality Watch Duration / User in online A/B tests, fully demonstrating its strong practical effectiveness in industrial production setting.
2. Related Work
Feature Interaction in Recommender Systems. Recommender systems often rely on rich, multi-domain feature inputs, and modeling interactions among these features is critical for accurate prediction. In early-stage systems, hand-crafted cross features were commonly used to explicitly capture such interactions (He et al., 2014). Factorization Machines (FM) (Rendle, 2010) emerged as a classical approach for modeling pairwise feature interactions, though they lack the capacity to model higher-order relationships. With the advent of Deep Neural Networks (DNNs), researchers observed that DNNs could implicitly model high-order feature interactions. This inspired hybrid approaches that combine both explicit and implicit modeling. Deep&Wide (Cheng et al., 2016) combines a linear component with DNNs, while DeepFM (Guo et al., 2017) and DLRM (Naumov et al., 2019) integrate FM into deep architectures. Models like xDeepFM (Lian et al., 2018) and DCN (Wang et al., 2017, 2021) introduce enhanced cross-networks that capture higher-order explicit interactions.
Attention-based Feature Interaction Architectures. With the success of attention mechanisms across multiple domains, attention-based models have gained traction for feature interaction modeling. AutoInt (Song et al., 2019) employs multi-head self-attention layers, while models such as InterHAt (Li et al., 2020) and ACN (Li et al., 2021) extend this design with task-specific enhancements. By transforming input features into unified tokens, attention layers can be applied effectively (Xu et al., 2023; Zhang et al., 2023; Gui et al., 2023). However, attention mechanisms were originally designed for natural language, where tokens are relatively independent and semantically clear. In contrast, feature tokens in recommendation systems are often context-dependent and semantically entangled. Hiformer (Gui et al., 2023) addresses this gap by introducing a heterogeneous attention layer. Yet, several open problems remain, such as mitigating token homogeneity in deeper layers and enabling explicit information exchange between tokens.
Scaling Laws in Ranking Models. Findings from large language models (LLMs) suggest that increasing model scale significantly improves performance (Kaplan et al., 2020). This trend has also been observed in large-scale ranking models. Wukong (Zhang et al., 2023) empirically verifies that scaling laws apply to recommendation models as well. While expanding embedding sizes and increasing hidden layer dimensions are direct scaling approaches, works like DLRM (Naumov et al., 2019) and DHEN (Xu et al., 2023) demonstrate that stacking diverse architectural modules can also yield performance gains under computational constraints. The Mixture-of-Experts (MoE) framework (Ma et al., 2018), originally proposed for multi-task learning, has proven effective across various domains (Fedus et al., 2022a; Lepikhin et al., 2020; Riquelme et al., 2021; Zhang et al., 2024; Liu et al., 2025). MoE models increase capacity by introducing multiple expert subnetworks, among which only a subset is activated during inference. This sparse activation strategy, known as Sparse MoE (Fedus et al., 2022a; Du et al., 2022), allows efficient training and inference while scaling models to billions of parameters, making it a promising technique for scaling large ranking models that require strict inference latency. Despite these advancements, there remains a lack of systematic analysis on the effectiveness of different scaling strategies, particularly in combining modular stacking and MoE architectures. While Wukong (Zhang et al., 2023) explores the impact of model depth and hidden dimension size under fixed FLOPs or parameter budgets, it focuses primarily on straightforward scaling methods—such as increasing layer count or width—and does not consider more advanced compositions like different token interaction strategies and sparse activations based on MoE.
Generative Recommenders. Another line of work focuses on Generative Recommenders (GRs) (Rajput et al., 2023; Zhai et al., 2024; Deng et al., 2025; Huang et al., 2025; He et al., 2025; Zheng et al., 2025). Unlike traditional recommender systems that predict user–item interaction scores, GRs directly generate future items based on the user behavior sequence. Despite recent advancements, GRs are unlikely to fully replace traditional approaches. In practice, GRs face significant challenges in training efficiency and inference latency due to the generative modeling process (Hou et al., 2025). Moreover, conventional recommender models can complement GRs by serving as reward models to guide generation (Deng et al., 2025). Therefore, GRs represent a parallel research direction and are beyond the scope of this work.
3. Method
3.1. Design Overview
As shown in Figure 1, our proposed ranking model architecture, Zenith (Optimized Efficient Neural Interaction with Tokenwise Handling), consists of a tokenization layer followed by stacked neural network layers. The tokenization layer processes both sparse categorical and dense features, projecting them into a compact set of high-dimensional representations, which we refer to as Prime Tokens. Each neural network layer then processes these Prime Tokens, and comprises two core components: the Token Fusion (TF) module and the Token Boost (TB) module.
Figure 1. High-level structure of the proposed Zenith architecture. Zenith employs a stack of layers, and each layer has a Token Fusion module and a Token Boost module. Detailed instantiations of these modules are shown in Figure 2.
The Token Fusion module captures feature interactions and serves a role analogous to traditional feature crossing. However, unlike explicit methods such as Factorization Machines (FM) or Cross Networks, the Token Fusion network is tailored to handling Prime Tokens. It performs both explicit inter-token crossing and implicit intra-token feature fusion, enabling the model to selectively focus on the most relevant interactions during training.
Next, the Token Boost module applies tokenwise transformations to enhance token heterogeneity and improve overall model expressiveness. As detailed in Section 4.3, maintaining token heterogeneity is critical when scaling up model depth. While conventional fully connected feed-forward networks tend to homogenize token representations as layers deepen, our proposed Token Fusion module mitigates this by encouraging each token to specialize in distinct information, thereby preserving heterogeneity across layers.
In the following subsections, we will first introduce the prime tokenization of features, followed by Token Fusion and Token Boost modules. Those module components can be instantiated with various architectural choices, and we will introduce two effective designs for each component, collectively formulating Zenith and Zenith++ versions of our architecture. While Zenith++ offers superior scalability and performance at larger model sizes, Zenith is simpler to implement and demands less hyperparameter tuning while also outperforming existing baselines. We will also discuss additional training and inference optimizations used in practice.
(a) Designs of the Token Fusion module. We use Retokenized Self-attention (RSA) for Zenith, and Tokenwise Multi-head Self-attention (TMHSA) for Zenith++. For Zenith, to compensate for the reduced number of output tokens due to retokenization, we employ an auxiliary MLP to generate additional tokens from input tokens.
(b) Designs of the Token Boost module. We use Tokenwise SwiGLU (TSwiGLU) for Zenith, and Tokenwise Sparse MoE (TSMoE) for Zenith++. For Zenith++, as shown on the right side, we illustrate the processing pipeline for a single token, where 2 out of 3 sparse experts are activated.
Figure 2. Designs of Token Fusion and Token Boost modules for Zenith and Zenith++.
3.2. Prime Tokenization of Features
Prior work (Zhang et al., 2023; Xu et al., 2023) has demonstrated the effectiveness of feature tokenization. For our livestreaming recommendation setting, we found that a smaller number of higher-dimensional Prime Tokens work better than using one or more tokens for each input feature. Consequently, after standard feature bucketization and embedding lookups, feature embeddings are mapped into a global token space using multi-layer perceptrons (MLPs). Features of similar semantics (e.g., user demographic features) are concatenated first and then mapped to a single prime token. Only important features, like user ID and sequence features, are mapped to separate tokens. As a result, $K$ input features would be mapped $T$ tokens, where $T\ll K$. Unlike traditional methods which rely on explicit feature interactions, the Token Fusion module automatically learns to filter and aggregate relevant token interactions during training. As a result, this approach significantly reduces resource consumption compared to traditional models, which often compute explicit interactions for all feature pairs. Moreover, this enables us to use more complex tokenwise methods to boost understanding for each Prime Token. Our detailed tokenization strategies are as follows:
  1. (1)
    ID features are assigned separate tokens. Embeddings of ID-type features (e.g., user ID, item ID) are critical for most interaction pairs and thus deserve explicit token representation.
  2. (2)
    Tokens should preserve complete feature embeddings. A single feature embedding should not be divided across multiple tokens. Preserving the integrity of feature embeddings is essential for effective token fusion.
  3. (3)
    Remaining tokens should be grouped together based on semantics with balanced information load. For non-ID features, we group embeddings such that each token contains a similar number of semantically similar feature embeddings, ensuring balanced information distribution across tokens.
3.3. Token Fusion
The Token Fusion module functions similarly to conventional feature interaction mechanisms, with the key difference that each Prime Token in our design may aggregate information from multiple features. 2(a) illustrates the Token Fusion design used in Zenith, as well as the enhanced version adopted in Zenith++.
Retokenized Self-Attention (RSA)
In Zenith, we employ a Retokenized Self-Attention network, which consists of a self-attention module followed by a retokenization component to facilitate token interaction (2(a), left). The self-attention mechanism extends the traditional FM structure by introducing a value component and a learnable weight matrix, enabling the model to focus on different parts of the input. Due to our Prime Token design, each token may encode multiple features that require further interaction. To address this, we introduce the retokenization module, where Prime Tokens exchange information, allowing features originally embedded within the same token to be fused with one another for subsequent interactions in the next self-attention layer.
The input Prime Tokens is a matrix $X\in\mathbb{R}^{T\times D}$, where $T$ is the number of tokens and $D$ is the token dimension. We compute the self-attention interaction via:
$O_{1}=XX^{\top}XW_{R},\quad W_{R}\in\mathbb{R}^{D\times k},\quad O_{1}\in\mathbb{R}^{D\times k}$
We then flatten and retokenize $O_{1}$ into $O_{TF}\in\mathbb{R}^{\hat{T}\times d}$ such that $Tk=\hat{T}D$. This retokenization is a computation-free operation that increases each token’s contextual richness, and enables the crossing of features within the same Prime Token.
Since $O_{TF}$ outputs $\hat{T}$ tokens, we use a MLP to add original input $X$ to $O_{TF}$ to build the residual network, with the final output $X_{TB}$:
$X_{TB}=\text{Norm}(O_{TF}+\text{MLP}(X))$
Tokenwise Multi-Head Self-Attention (TMHSA)
For Zenith++, we upgrade the Fused Self-Attention network to Tokenwise Multi-Head Attention for token fusion. TMHSA significantly enhances model expressiveness by assigning distinct query, key, and value projection matrices to each Prime Token, thereby allowing the model to host substantially more parameters. Unlike traditional attention mechanisms that share projection weights across tokens, tokenwise attention introduces unique projections per token, enabling richer interactions without a substantial increase in computational cost. As TMHSA provides sufficient capacity for Prime Tokens to fully interact and exchange information, we remove the retokenization module in Zenith++.
Specifically, denoting the $i$-th token as $t_{i}$, the output of the multi-head attention layer is computed as:
$O_{TF}=\text{concat}\left(\left\{\frac{Q_{h}K_{h}^{\top}}{\sqrt{d_{k}}}V_{h}\right\}_{h=1}^{H}\right)$
where
$Q_{h}=\text{concat}\left(\{t_{i}q_{(i,h)}\}_{i=1}^{n}\right)$
$K_{h}=\text{concat}\left(\{t_{i}k_{(i,h)}\}_{i=1}^{n}\right)$
$V_{h}=\text{concat}\left(\{t_{i}v_{(i,h)}\}_{i=1}^{n}\right)$
Here, $H$ is the number of attention heads, and $d_{k}=d/H$ is the dimensionality of each head. The projection weights $q_{(i,h)},k_{(i,h)},v_{(i,h)}\in\mathbb{R}^{d\times d_{k}}$ are token- and head-specific. The input of Token Boost $X_{TB}$ after residual is:
$X_{TB}=\text{Norm}(O_{TF}+X)$
3.4. Token Boost
The Token Boost module plays a crucial role in enhancing model expressiveness and enhancing token heterogeneity, which, as we will demonstrate later in Section 4.3, is key to efficient scaling up. The Prime Token design enables each token to capture rich and diverse information, effectively functioning as a collection of specialized experts, akin to those in a Mixture-of-Experts network. To complement it, our Token Boost module design is centered around tokenwise parameterizations, where each token has its own token boost network.
Tokenwise SwiGLU (TSwiGLU)
For the Zenith version of Token Boost, we use Tokenwise SwiGLU (2(b) left):
$O_{TB}=\text{SwiGLU}(X_{TB})=(\text{Swish}(X_{TB}W_{1})\otimes X_{TB}W_{2})W_{3}$
where $\text{Swish}(t)=t\cdot\text{sigmoid}(t)$, $W_{1},W_{2}\in\mathbb{R}^{D\times r}$, $W_{3}\in\mathbb{R}^{r\times D}$, and $\otimes$ denotes the elementwise (Hadamard) product. This elementwise product increases the order of multiplicative interactions similar to those in the DCN-V2 network (Wang et al., 2021), which motivates our choice of SwiGLU as the backbone for the feedforward neural network.
Since the $X_{TB}$ of Zenith only consists of $\hat{T}$ tokens, we regenerate the remaining $T-\hat{T}$ tokens via a lightweight MLP applied to $X$. The final output of Zenith is:
$O_{Zenith}=\text{Norm}(\text{concat}(O_{TB},\text{MLP}(X))+X)$
Tokenwise Sparse Mixture-of-Experts (TSMoE)
To further improve the scaling efficiency of the model, Zenith++ replaces the standard feedforward neural network with a Tokenwise Sparse Mixture-of-Experts, inspired from the standard Sparse Mixture-of-Experts (Fedus et al., 2022a; Du et al., 2022). As illustrated in 2(b) right, this module comprises two types of experts: shared experts, which are always involved in the computation, and sparse experts, which are selectively activated based on token-specific routing decisions. This design enables significantly greater model capacity without additional computational overhead.
Let $E_{c}$ be the number of shared (common) experts, $E_{s}$ the total number of sparse experts, and $E_{a}$ the number of sparse experts activated per token. For each token $t_{i}$, a gating mechanism determines which experts to activate:
$\text{Gate}(t_{i})=W_{0}^{i}t_{i}$
where $W_{0}^{i}\in\mathbb{R}^{D\times E_{s}}$ is the routing weight matrix specific to token $t_{i}$. The top-$E_{a}$ elements of $\text{Gate}(t_{i})$ determine the activated expert indices.
The output of the MoE-enhanced FFN for token $t_{i}$ is:
$\text{TB}(t_{i})=\sum_{e_{c}=1}^{E_{c}}\text{SwiGLU}^{e_{c}}(t_{i})+\sum_{e_{a}\in\text{TopK}(\text{Gate}(t_{i}),E_{a})}\text{SwiGLU}^{e_{a}}(t_{i})$
Here, $\text{SwiGLU}^{e_{c}}$ and $\text{SwiGLU}^{e_{a}}$ denote the $e_{c}$-th shared expert and $e_{a}$-th activated sparse expert, respectively, each implemented using the TSwiGLU function, which is the same as Zenith. After each token goes through Sparse MoE, we obtain the output of Token Boost as $O_{TB}$.
Finally, the module output is passed through a residual connection followed by Layer Normalization:
$O_{Zenith}=\text{Norm}(O_{TB}+X)$
3.5. Training and Inference Optimizations
Zenith hinges on fused‐token representations and token-wise computations. While these tokenwise operations improve modeling flexibility, they also inflate the total number of matrix multiplications, reducing throughput on modern accelerators. Introducing a Sparse-MoE layer adds two further complications: (i) expert-activation imbalance, where a minority of experts receive most traffic, and (ii) the overhead of routing computation solely to the selected experts. We tackle these issues with several complementary strategies, detailed in the following sections.
Efficient Tokenwise Computations.
To eliminate the kernel-launch overhead that accrues when each token triggers an independent matrix multiplication, we adopt the GroupedGEMM primitive available in NVIDIA’s cuBLAS library (Hejazi, 2024). GroupedGEMM packs the $N_{\text{tok}}$ token-wise GEMM operations into a single batched call, yielding markedly higher GPU occupancy and end-to-end throughput for our tokenwise designs.
Learning-Rate Warm-up.
Early in training, the Sparse-MoE router tends to channel most tokens through only one or two experts, depriving the others of gradient signals. We counteract this collapse with an extended learning-rate warm-up (LR warmup): training starts at just 0.1% of the base learning rate and increases linearly to the nominal value over the first one million steps. The resulting low-gradient regime gives the randomly initialized router time to explore diverse routing patterns, after which expert loads become well balanced and every expert receives sufficient training data.
Auxiliary Losses.
To further encourage balanced workload among experts, we introduce two auxiliary loss terms. The first is the load balancing loss (Fedus et al., 2022b), which penalizes the mismatch between the router’s output probabilities and the actual number of tokens routed to each expert:
$\mathcal{L}_{\text{load}}=\alpha\cdot\frac{1}{BTE_{s}}\sum_{i=1}^{E_{s}}\left(\frac{1}{E_{a}}f_{i}\right)\bar{\pi}_{i}$
where
$f_{i}=\frac{1}{TB}\sum_{b=1}^{B}\sum_{t=1}^{T}m_{b,t,i}$
is the fraction of tokens actually routed to expert $i$, and
$\bar{\pi}_{i}=\frac{1}{TB}\sum_{b=1}^{B}\sum_{t=1}^{T}\pi_{b,t,i}$
is the average routing probability assigned to expert $i$.
To further prevent expert collapse (i.e., one expert dominating the routing decisions), we employed a second term—referred to as the z-loss (Zoph et al., 2022)—to discourage the router from producing overly large logits:
$\mathcal{L}_{z}=\beta\cdot\frac{1}{BT}\sum_{b=1}^{B}\sum_{t=1}^{T}\left[\ln\left(\sum_{i=1}^{E_{s}}e^{z^{b,t,i}}\right)\right]^{2}$
where $z^{b,t,i}$ denotes the router’s pre-softmax output for expert $i$ at position $(b,t)$.
4. Evaluation
We evaluate the proposed Zenith architecture (with Zenith and Zenith++ implementations) against other popular architectures across different model scales. To demonstrate the practical effectiveness of our method, all models are deployed on the TikTok Live recommendation platform.
4.1. Evaluation Setup
4.1.1. Dataset
We leverage the production dataset from the ranking stage of the TikTok Live recommendation platform for training and evaluation, which contains real-world user-item interaction data in a streaming fashion. The statistics of the dataset is shown in Table 1.
Table 1. Statistics of the dataset.
# Instances # Features # Targets
TikTok Live 168B 4552 98
4.1.2. Metrics
We adopt multiple metrics to comprehensively assess both the computational cost and predictive performance of each model:
  • Model Scale. We quantify model size and compute overhead with:
    • #Params: Total number of parameters in the model, including trainable weights from all layers. The sparse embedding table is shared across models in our experiments and excluded from comparison.
    • GFLOP: Total number of floating-point operations (in billions) required for a single inference. This provides a more accurate measure of computational load, especially when not all parameters are actively used.
  • Model Performance: We report the following standard model performance metrics for our main CTR target:
    • LogLoss: The logarithmic loss of the predicted probabilities against the ground truth.
    • AUC: Area Under the ROC Curve, measuring the model’s ability to rank positive samples above negatives.
    • UAUC: User-Averaged AUC, which averages AUC scores per user to reflect personalized ranking quality.
All results are reported as both raw values and relative improvements over our baseline model in production.
4.1.3. Baselines & Hyperparameters
We chose some widely used DLRMs as well as some recent strong baselines, including DCN-V2 (Wang et al., 2021), DHEN (Xu et al., 2023), and Wukong (Zhang et al., 2023). We tuned the hyperparameters to achieve the scaling we need for the baselines as follows:
  • DCN-V2: We scaled up the model by increasing the number of layers, as well as using an MLP in the Cross Network instead of a single projection matrix. We applied sequential modeling with a Deep Network after the Cross Network.
  • DHEN: We used a combination of DCN-V2’s cross network and a multi-head self-attention module. We scaled up the model by increasing the number of DHEN layers and the size of the MLPs attached after each interaction module.
  • Wukong: We scaled up the model by increasing the number of Wukong layers and increasing the number of hidden units of the MLP in its FM Block.
For Zenith and Zenith++, we employ 32 Prime Tokens grouped from the original feature embeddings, including 3 tokens based on ID features, 13 tokens based on sequence features, and 16 other tokens based on other generalizable features. The dimension of each Prime Token ranges from 256 to 1024, and the number of layers range from 2 to 8, depending on the model scale.
4.1.4. Training
For all models, we use the RMSPropV2 optimizer with a learning rate of 0.01, momentum of 0.99999, and an initialization factor of 0.015625, along with learning rate warm-up.
4.2. Offline Results
(a) Relative LogLoss vs. Model Size.
(b) Relative LogLoss vs. Compute.
Figure 3. Model performance comparison between our proposed architectures and the baselines. Note that in (a), for Zenith++, we also report the effective parameter count per inference—denoted Zenith++(act)—which reflects the reduced active parameters enabled by our Tokenwise Sparse MoE. All LogLoss values are shown against the baseline DCN-V2 model.
Table 2. Model Performance at Different Scales. Values in parentheses indicate relative improvement over the Baseline DCN-V2 model. For Zenith++, since we use Sparse MoE, we report the activated parameter count/total parameter count.
Scale Architecture #Params (M) GFLOP (per example) AUC Logloss UAUC
Baseline DCN-V2 22.8 0.0302 .81151 .08255 .69627
Small DCN-V2 98.5 0.129 .81290 (+0.17%) .08237 (-0.22%) .69871 (+0.35%)
DHEN 136 0.140 .81351 (+0.25%) .08227 (-0.34%) .69877 (+0.36%)
Wukong 105 0.143 .81221 (+0.09%) .08247 (-0.10%) .69688 (+0.09%)
Zenith 106 0.0800 .81404 (+0.31%) .08220 (-0.42%) .69999 (+0.53%)
Zenith++ 90/150 0.0494 .81438 (+0.35%) .08214 (-0.50%) .70036 (+0.59%)
Medium DCN-V2 522 0.686 .81283 (+0.16%) .08237 (-0.22%) .69881 (+0.36%)
DHEN 452 0.440 .81409 (+0.32%) .08219 (-0.44%) .69990 (+0.52%)
Wukong 493 0.646 .81287 (+0.17%) .08238 (-0.21%) .69822 (+0.28%)
Zenith 468 0.250 .81479 (+0.40%) .08208 (-0.57%) .70085 (+0.66%)
Zenith++ 383/916 0.202 .81521 (+0.46%) .08202 (-0.64%) .70170 (+0.78%)
Large DCN-V2 1050 1.37 .81299 (+0.18%) .08235 (-0.24%) .69876 (+0.36%)
DHEN 952 0.906 .81448 (+0.37%) .08213 (-0.51%) .70036 (+0.59%)
Wukong 946 1.25 .81306 (+0.19%) .08234 (-0.25%) .69850 (+0.32%)
Zenith 928 0.519 .81509 (+0.44%) .08203 (-0.63%) .70148 (+0.75%)
Zenith++ 920/1822 0.698 .81594 (+0.55%) .08188 (-0.81%) .70303 (+0.97%)
(a) Zenith w/o TSwiGLU
(b) Zenith with TSwiGLU
(c) Zenith++ w/o TSMoE
(d) Zenith++ with TSMoE
Figure 4. The comparison of token similarity with and without tokenwise operations in the Token Boost network for Zenith (TSwiGLU) and Zenith++ (TSMoE). The similarity level is defined by the absolute cosine similarity of the token pair, and the deeper red color indicates more similar output, which is detrimental for the deeper layer’s expressiveness, limiting model scaling capability. We sampled the first 10 tokens for clarity, and used the average token similarity of the output tokens from the third layer in each model.
Figure 3 presents a comparative analysis of relative LogLoss across various model sizes and compute budgets. We report results from two perspectives: (a) performance versus model size, and (b) performance versus inference-time FLOPs. The baseline is a 22.8M-parameter DCN-V2 model. Although DCN exhibits moderate gains when scaled to 100M parameters, further increasing the number of DCN layers or expanding its MLP does not yield additional improvements. In contrast, Wukong scales more effectively than DCN but only outperforms it at scales of 500M parameters or more, and the performance gain remains modest. Wukong employs a large MLP to map interaction results back to tokens, which incurs high parameter costs and leads to inefficient parameter utilization—limiting its advantage to very large scales.
DHEN integrates DCN-v2’s cross network with self-attention, delivering consistently stronger performance than both DCN and Wukong across all scales, with LogLoss improvements of approximately 0.3%–0.5%. DHEN also shows good scalability. However, its benefit remains marginal, and its hybrid architecture increases tuning complexity due to the heterogeneous components involved.
Our proposed Zenith model consistently outperforms all baselines across different scales, achieving 0.42%–0.63% lower LogLoss compared to the DCN baseline. With our efficient design, including the use of Prime Tokens to reduce input token count and retokenization for lightweight token fusion, Zenith achieves better performance while minimizing training FLOPs. When compared to the best baseline, DHEN, at equivalent compute levels, Zenith improves LogLoss by 0.15%–0.2%, which represents a significant gain in real-world scenarios. Additionally, Zenith’s simpler and more modular architecture makes operator-level optimizations and tuning easier in production environments.
Zenith++ further enhances model performance by incorporating TMHSA and TSMoE. These components introduce sparsity and per-token computation to maintain expressiveness while reducing computational cost. As a result, Zenith++ surpasses Zenith in performance while keeping overall FLOPs under control. On average, Zenith++ achieves a 0.08% lower LogLoss than Zenith when activating a similar number of parameters. Notably, at scales of 500M parameters or fewer, Zenith++ actually uses fewer FLOPs. Furthermore, these curves indicate that Zenith++ holds even greater potential for further scaling.
As shown in Table 2, Zenith and Zenith++ consistently achieve the highest AUC and UAUC across all model sizes. In the small-scale setting, Zenith significantly outperforms DCN-V2, Wukong, and DHEN. At medium scale, Zenith++—with only 383M active parameters out of 916M total—achieves a 0.78% UAUC gain over the DCN baseline, outperforming all other models. Given its strong performance and efficient computation, we adopted this version of Zenith++ for real-world online A/B testing. At the large scale, Zenith++ (with 920M active out of 1822M total parameters) delivers the best results, reaching an AUC gain of 0.55% and a UAUC gain of 0.97%, establishing Zenith++ as the state-of-the-art architecture.
Overall, Zenith models demonstrate stronger scalability and higher efficiency. Across all scales, Zenith not only reduces prediction loss but also offers favorable compute-performance trade-offs, making it an ideal architecture for real-world large-scale recommender systems such as TikTok Live.
4.3. Token Heterogeneity
We now highlight the key advantage of our tokenwise parameterization scheme in the Token Boost module, which enabled superior scaling laws for both Zenith and Zenith++.
Initially, we attempted to apply traditional non-tokenwise SwiGLU and SMoE in our Token Boost modules, but encountered significant performance saturation after 50M parameter scale. We tried to increase the Prime Token dimension from 512 to 768, increase the number of tokens from 8 to 16, and increase the number of layers from 4 to 6, all of which doubled the total number of parameters without any performance gains. These observations motivated us to deep dive into our architecture and understand how our Prime Tokens interact with one another in the model. Surprisingly, as shown in 4(a) and 4(c), we found that the Prime Tokens become almost homogeneous after feature interaction by a few layers. This indicates that the models can no longer extract additional information from the feature tokens by stacking more interaction layers, explaining the performance saturation.
To address this, we need to ensure that the Prime Tokens remain heterogeneous after each interaction layer, carrying distinctive information for further extraction. We achieve this by upgrading the traditional SwiGLU and SMoE modules with tokenwise parameterization. This provides the needed model capacity to allow each Prime Token to capture different information, leading to strong token heterogeneity even after many layers of feature interaction. As demonstrated in 4(b) and 4(d), with tokenwise parameterizations, the Token Boost modules significantly reduces the average absolute similarity—from 0.5 to 0.06 in Zenith, and from 0.68 to 0.47 in Zenith++, allowing both variants to further scale up without reaching performance plateaus.
4.4. Ablation Study
We perform an ablation study to evaluate the impact of key design choices in Zenith++. As summarized in Table 3, both TSMoE and TMHSA, contribute substantially to model performance, while training techniques such as learning rate warmup and auxiliary loss provide additional gains. This confirms that each component plays a positive role, with architectural designs yielding the most significant improvements. Meanwhile, substituting prime tokenization with an MLP-based auto-grouping approach incurs only a 0.09% AUC degradation, demonstrating that Zenith is robust even without expert-tuned feature tokenization.
Table 3. Ablation results of Zenith++.
Ablated Component AUC ($\Delta$%)
Disable Token Fusion 0.60-0.60
Disable Token Boost 0.32-0.32
Replace TSMoE with MoE 0.10-0.10
Replace TMHSA with Self-Attention 0.12-0.12
Remove LR Warmup 0.04-0.04
Remove Auxiliary Loss 0.01-0.01
Replace Prime Tokenization with Auto-grouping 0.09-0.09
4.5. Online A/B Results
To fully validate the effectiveness of the proposed architecture in real-world scenarios, we conducted a month-long online A/B test comparing the DCN-V2 baseline and our Zenith++ (383/916M) model for TikTok Live, a leading online livestreaming platform that attracts billion of users globally. We present the online performance metrics—AUC and Logloss, as well as two sets of key business indicator metrics for livestreaming recommendation—Quality Watch Session / User and Quality Watch Duration / User, where the Quality Watch Session is defined as a watch session that is either deep (e.g., greater than 1 min) or accompanied with positive user actions (e.g., like, comment, follow, share, gift, etc), and the Quality Watch Duration is defined as the total time a user spend in their Quality Watch Sessions. The results are shown in Table 4. Zenith significantly outperforms the baseline in both online AUC and logloss, and realizing gains of +9.93% in Quality Watch Session / User and +8.11% in Quality Watch Duration / User. These results highligh the substantial real-world impacts of our novel ranking architecture.
Table 4. Online performance gains across key metrics.
Online Metrics Performance Gains
CTR AUC +1.05%
CTR Logloss -1.10%
Quality Watch Session / User +9.93%
Quality Watch Duration / User +8.11%
5. Conclusion
In this work, we present Zenith, a scalable and efficient neural architecture tailored for large-scale recommendation systems. By introducing the concept of Prime Tokens and designing dedicated Token Fusion and Token Boost modules, Zenith effectively balances model expressiveness and computational efficiency, exhibiting superior scaling laws compared to existing approaches. We also highlight the importance of token heterogeneity in scaling up ranking models, and demonstrated practical ways to promote it in our architectural designs. We hope that our work will be an important step towards optimized user experience through personalized recommendation.
Appendix A Complexity Analysis
In this section, we present a formal complexity analysis of our proposed Zenith and Zenith++ models.
Zenith:
Let $T$ be the number of tokens, $D$ be the token dimension, $k$ be the intermediate projection dimension, and $r$ be the hidden size of the SwiGLU. The self-attention interaction computes $XX^{\top}XW_{R}$ with $W_{R}\in\mathbb{R}^{D\times k}$, contributing $Dk$ parameters and incurring $O(T^{2}D+TDk)$ compute. The retokenization step reshapes $O_{1}\in\mathbb{R}^{T\times k}$ to $O_{TF}\in\mathbb{R}^{\hat{T}\times D}$ (where $Tk=\hat{T}d$) without computation or parameters.
The Token Boost module applies a SwiGLU activation with weights $W_{1},W_{2}\in\mathbb{R}^{D\times r}$ and $W_{3}\in\mathbb{R}^{r\times D}$, totaling $3\hat{T}Dr$ parameters and $O(\hat{T}Dr)$ compute. A lightweight MLP processes the remaining $T-\hat{T}$ tokens, adding $D^{2}\frac{T-\hat{T}}{T}$ parameters and $O((T-\hat{T})D^{2})$ compute. The final output includes a residual connection and normalization with negligible cost.
In total, the model has $Dk+D^{2}+3\hat{T}Dr$ parameters and a computation complexity of $O(T^{2}D+TDk+\hat{T}Dr+(T-\hat{T})D^{2}$. Since $D,k,r$ are normally on the same scale, and that $\hat{T},T\ll D,k,r$, allowing them to be treated as constant values, the computation complexity is close to $O(D^{2})$.
Zenith++:
In tokenwise attention, each of the $n$ tokens has unique projection matrices for each of the $H$ heads, leading to a total parameter count of $3THDd_{k}=3TD^{2}$ (since $d_{k}=D/H$). The computation cost consists of tokenwise projections ($O(TD^{2})$) and attention computation ($O(T^{2}D)$), yielding an overall complexity of $O(TD^{2}+T^{2}D)$.
For the sparse MoE layer, the routing weights contribute $TDE_{s}$ parameters, and all $E_{c}$ shared and top-$E_{a}$ sparse experts contribute a total of $3Dr(E_{c}+E_{s})$ parameters. Each token’s routing and expert computation require $O(DE_{s}+(E_{c}+E_{a})dr)$ operations, resulting in total compute $O(TDE_{s}+n(E_{c}+E_{a})Dr)$.
Since $T$ can be treated as a constant, $E_{c}$, $E_{a}$, and $E_{s}$ are constants, and $D,r$ are on the same scale, the parameter count simplifies to $O(D^{2})$, and the compute complexity is close to $O(D^{2})$ as well.